Please show all work, you to avoid using excel
Chapter 10 Exercise Problem McGee Computing Service is considering an average-risk project that will cost $9 and is expected to generate annual after-tax cash inflow in the amount of $19,000 for the next 12 years. The firm has 300,000 shares of common stock outstanding at a price of $40 a share, 10,000 shares of preferred stock outstanding at a price of $80 a share, and 2,000 bonds outstanding at a price of 77.40 percent of par. The companys 9% preferred stock has a par value of $100 each. Its currently outstanding bonds have a face value of $1,000, carry an annual coupon rate of 8%, and will mature in 10 years. McGees most recently paid common dividends are $4 a share and the expected growth in dividends is 6% per year indefinitely. It has no intention to issue additional common equity and instead will rely on retaining earnings for common equity financing. The companys applicable tax rate is 25%. (Assume yearly coupon payments on the bonds.)
Answer #1

Price of stock = D1/ (Ke- g)
where D1 = dividend to be paid next year = 4 + 6% = $4.24
price of stock = $40
g = growth = 6%
hence,
40 = 4.24/ (ke - 6%)
solving the above equation, we get, ke = 10.6%

hence, cost of common stock = ke = 10.6%
cost of preferred stock = kp = 9%
cost of bond = kd = 8% - 25% tax
= 8%*(1-0.25)
= 6%

Market value of common stock = 300,000*40 = $12,000,000
Market value of preferred stock = 10,000*80 = $800,000
Market value of bond = 2000*77.40 = $154,800
Total market value = $12,954,800

Computing WACC using market value weights:
WACC for first 10 years (until bonds mature)
= {(10.6%*12,000,000) + (9%*800,000) + (6%*154,800)} /12,954,800
= 1,353,288/ 12,954,800
= 10.45%

WACC for last 2 years (no bonds outstanding)
= {(10.6%*12,000,000) + (9%*800,000)} /12,800,000
= 1,344,000/ 12,800,000
= 10.5%

PV of after tax cash inflows discounted at WACC
= 19,000 at 10.45% for 10 years + 19,000 at 10.5% for last 2 years
= (19,000* 6.0275) + (19,000* 0.6352)
= $126,591

Hence NPV of the proposed project at time zero = 126,591 - 92,000
= $ 34,591
Since the project has a positive NPV it is suggested to take up the project.